# YITP-SB-2025-20

### v1

@@ 3 $SU(2)$ $N_{f}=4$ series
In this section we discuss the case of $SU(2)$ with four flavors. It was observed in [Deb:2025ypl], that for certain rational values of $\alpha$, the generalized Schur partition function coincides with the Schur indices of theories in rank-one Deligne-Cvitanović (DC) series. Theories belonging to this series are labelled by their flavor symmetry $\mathfrak{g}$ belonging to the Deligne-Cvitanović series of simple Lie algebras
@@
The theories are also known as $\mathfrak{a}_{0}\simeq(A_{1},A_{2})$, $\mathfrak{a}_{1}\simeq(A_{1},A_{3})$, $\mathfrak{a}_{2}\simeq(A_{1},D_{4})$ Argyes-Douglas theories [Argyres:1995jj, Argyres:1995xn] and the $\mathfrak{e}_{6}$$\mathfrak{e}_{7}$$\mathfrak{e}_{8}$ are known as Minahan-Nemeschansky theories [Minahan:1996fg, Minahan:1996cj]. $\mathfrak{d}_{4}$ is the only Lagrangian theory in this series and corresponds to $SU(2)$ theory with four flavors. $\mathfrak{f}_{4}$ and $\mathfrak{g}_{2}$ do not correspond to known 4d theories. We will denote the generalized Schur partition function of these theories by $\hat{\mathcal{Z}}_{\mathfrak{g}}(q,\alpha)$. The generalized Schur partition function for the $\mathfrak{d}_{4}$ theory takes the following form
@@
Now, we can use the above $q$-series and evaluate it for any value of $\alpha$. As a series expansion in $q$, the above coincides with the following expression as a power series in $q$ [Franc_Mason_2014]
@@
Table 3 collects the values of $\alpha$ for which we obtain $q$-series with integer coefficients. Based on the $q$-series, we note down the VOA to which these vacuum characters correspond to. These coincide with the VOA characters found in [Kim:2024dxu] using trace of higher powers of the quantum monodromy. To make the connection precise, for the theory with flavor symmetry $\mathfrak{g}$, we rescale $\alpha$ as follows
@@
In this rescaling $\alpha_{\mathfrak{g}}=1$ corresponds to the generalized Schur partition function for the theory $\mathfrak{g}$. For the case of Argyres-Douglas theories in this series, this limit coincides with generalized Schur limit for the respective theory using $\mathcal{N}=1$ Lagrangian description of [Maruyoshi:2016tqk, Maruyoshi:2016aim, Agarwal:2016pjo], with the overall rescaling of $\alpha$. For $\mathfrak{g}=\mathfrak{a}_{0},\mathfrak{a}_{\frac{1}{2}},\mathfrak{a}_{2},\mathfrak{g}_{2}$
@@
where ${\chi}_{\mathfrak{g}}$ denotes the character in terms of fugacities which reproduce the correct flavored index  [^9] The embedding of the flavor fugacities in $SO(8)$ can be inferred from the Lagrangians in [Maruyoshi:2016tqk, Maruyoshi:2016aim, Agarwal:2016pjo].
@@ 4.1 $USp(2N)$ $N_{f}=2N+2$ series
The values of $\alpha$ for which we find integer coefficient $q$-series are collected in table 5 and 6. For the $USp(4)$ case, we find that for $\alpha=-\frac{1}{7},-\frac{2}{7}$ correspond to the $q$-series expansion of $\mathfrak{osp}(1|4)_{1}$ and $(X_{1})_{1}$ VOA. $\alpha=-\frac{1}{5}$ corresponds to the $q$-series expansion of $\left(\mathfrak{a}_{4}\right)_{1}$. As discussed in previous sections, by an overall rescaling of $\alpha$, these results can be interpreted as generalized Schur limits of the Argyres-Douglas theories and are equal to the trace of higher powers of the quantum monodromy considered in [Kim:2024dxu], which is consistent with our conjecture. We refer the reader to [Kim:2024dxu] for the computation of these traces. The $USp(6)$ case for $\alpha=-\frac{1}{7},-\frac{1}{9}$ correspond to the characters of $(\mathfrak{a}_{6})_{1}$ and $\mathfrak{osp}(1|6)$ [Creutzig:2024ljv]. We perform the flavored checks below.
@@ Appendix A Notations and MLDE coefficients
We follow the notations in [Beem:2017ooy] (see also [Beem:2021zvt]). Here we collect only relevant formulae and notation. A monic modular linear differential operator (MLDO) is defined as

### v2

@@ 3 $SU(2)$ $N_{f}=4$ series
In this section we discuss the case of $SU(2)$ with four flavors. It was observed in [21], that for certain rational values of $\alpha$, the generalized Schur partition function coincides with the Schur indices of theories in rank-one Deligne-Cvitanović (DC) series. Theories belonging to this series are labelled by their flavor symmetry $\mathfrak{g}$ belonging to the Deligne-Cvitanović series of simple Lie algebras
@@
The theories are also known as $\mathfrak{a}_{0}\simeq(A_{1},A_{2})$, $\mathfrak{a}_{1}\simeq(A_{1},A_{3})$, $\mathfrak{a}_{2}\simeq(A_{1},D_{4})$ Argyes-Douglas theories [3, 4] and the $\mathfrak{e}_{6}$$\mathfrak{e}_{7}$$\mathfrak{e}_{8}$ are known as Minahan-Nemeschansky theories [35, 36]. $\mathfrak{d}_{4}$ is the only Lagrangian theory in this series and corresponds to $SU(2)$ theory with four flavors. $\mathfrak{f}_{4}$ and $\mathfrak{g}_{2}$ do not correspond to known 4d theories. We will denote the generalized Schur partition function of these theories by $\hat{\mathcal{Z}}_{\mathfrak{g}}(q,\alpha)$. The generalized Schur partition function for the $\mathfrak{d}_{4}$ theory takes the following form
@@
Now, we can use the above $q$-series and evaluate it for any value of $\alpha$. As a series expansion in $q$, the above coincides with the following expression as a power series in $q$ [23]
@@
Table 3 collects the values of $\alpha$ for which we obtain $q$-series with integer coefficients. Based on the $q$-series, we note down the VOA to which these vacuum characters correspond to. These coincide with the VOA characters found in [28] using trace of higher powers of the quantum monodromy. To make the connection precise, for the theory with flavor symmetry $\mathfrak{g}$, we rescale $\alpha$ as follows
@@
In this rescaling $\alpha_{\mathfrak{g}}=1$ corresponds to the generalized Schur partition function for the theory $\mathfrak{g}$. For the case of Argyres-Douglas theories in this series, this limit coincides with generalized Schur limit for the respective theory using $\mathcal{N}=1$ Lagrangian description of [33, 32, 1], with the overall rescaling of $\alpha$. For $\mathfrak{g}=\mathfrak{a}_{0},\mathfrak{a}_{\frac{1}{2}},\mathfrak{a}_{2},\mathfrak{g}_{2}$
@@
where ${\chi}_{\mathfrak{g}}$ denotes the character in terms of fugacities which reproduce the correct flavored index  [^10] The embedding of the flavor fugacities in $SO(8)$ can be inferred from the Lagrangians in [33, 32, 1].
@@ 4.1 $USp(2N)$ $N_{f}=2N+2$ series
The values of $\alpha$ for which we find integer coefficient $q$-series are collected in table 5 and 6. For the $USp(4)$ case, we find that for $\alpha=-\frac{1}{7},-\frac{2}{7}$ correspond to the $q$-series expansion of $\mathfrak{osp}(1|4)_{1}$ and $(X_{1})_{1}$ VOA. $\alpha=-\frac{1}{5}$ corresponds to the $q$-series expansion of $\left(\mathfrak{a}_{4}\right)_{1}$. As discussed in previous sections, by an overall rescaling of $\alpha$, these results can be interpreted as generalized Schur limits of the Argyres-Douglas theories and are equal to the trace of higher powers of the quantum monodromy considered in [28], which is consistent with our conjecture. We refer the reader to [28] for the computation of these traces. The $USp(6)$ case for $\alpha=-\frac{1}{7},-\frac{1}{9}$ correspond to the characters of $(\mathfrak{a}_{6})_{1}$ and $\mathfrak{osp}(1|6)$ [19]. We perform the flavored checks below.
@@ Appendix A Notations and MLDE coefficients
We follow the notations in [6] (see also [7]). Here we collect only relevant formulae and notation. A monic modular linear differential operator (MLDO) is defined as